# Data Valuation using Neural Networks for Efficient Instruction Fine-Tuning

### v2

@@ Abstract
Influence functions provide crucial insights into model training, but existing methods suffer from large computational costs and limited generalization. Particularly, recent works have proposed various metrics and algorithms to calculate the influence of data using language models, which do not scale well with large models and datasets. This is because of the expensive forward and backward passes required for computation, substantial memory requirements to store large models, and poor generalization of influence estimates to new data. In this paper, we explore the use of small neural networks – which we refer to as the InfluenceNetwork – to estimate influence values, achieving up to 99% cost reduction. Our evaluation demonstrates that influence values can be estimated with models just 0.0027% the size of full language models (we use 7B and 8B versions). We apply our algorithm of estimating influence values (called NN-CIFT: Neural Networks for effiCient Instruction Fine-Tuning) to the downstream task of subset selection for general instruction fine-tuning. In our study, we include four state-of-the-art influence functions and show no compromise in performance, despite large speedups, between NN-CIFT and the original influence functions. We provide an in-depth hyperparameter analyses of NN-CIFT. The code for our method can be found here: https://github.com/agarwalishika/NN-CIFT.
Data Valuation using Neural Networks for Efficient Instruction Fine-Tuning
Ishika Agarwal UIUC ishikaa2@illinois.edu                        Dilek Hakkani-Tür UIUC dilek@illinois.edu
@@ 1 Introduction
Table 1: Approximate computational complexity of data valuation in previous works measured by the cost of forward passes ($F$) or the cost of backward passes ($B$) through a model. $M$ and $N$ are the cardinality of $\mathcal{D_{F}}$ and $\mathcal{D_{T}}$, a fine-tuning and target dataset respectively, we use for subset selection. See Appendix B.1 for more details. Size denotes the number of parameters of the corresponding model. Note: larger models have a higher cost for forward and back passes.
In this paper, we introduce NN-CIFT: Neural Networks for effiCient Instruction Fine-Tuning and explore how to train influence functions efficiently. We improve efficiency by using compact neural networks – which we coin as the InfluenceNetwork – that are 0.0027% the size of LLMs, to estimate influence. Figure 1 outlines our methodology with a pairwise influence function (more details about pairwise influence functions in Appendix B.1).
@@
    alleviates the cost of using expensive LLMs during data valuation by using smaller and cheaper neural networks, without affecting the performance on downstream tasks (Tables 2 and 3);
@@
Section 2 outlines the current state of research in data valuation and data selection. Section 3 explains the problem setting. Section 4 presents the main methodology for NN-CIFT and motivating results. Finally, Section 5 reports results on the downstream task of subset selection after the data valuation stage. In our evaluation, we find that using a small LLM with the original influence functions results in degraded performance. Our hyperparameter studies are in Section 4.4, Figure 3 and Section 5.2, Figure 4. Lastly, the SOTA influence functions are detailed in Appendix B.
@@ 3 Problem Formulation
As mentioned before, computing influence functions can be a very expensive process. There are two kinds of influence functions: pairwise and pointwise – both require forward/backward passes through language models, but the costs slightly differ. Pairwise influence functions compute the influence between every pair of points in a dataset. We study three SOTA pairwise functions, whose formulations are details in Appendix B.1. This paper also studies one pointwise influence functions that simply compute the influence of each data point individually, formally outlined in Appendix B.2. While pointwise influence functions are more efficient than pairwise, they are not as performant during subset selection Xia et al. (2024); Agarwal et al. (2025).
@@ 4 Learning Influence Estimation
$0\leq IN_{\theta}(i,j)$ $\leq 1,$
$0\leq\theta(\texttt{concat}(\texttt{emb}(i),\texttt{emb}(j)))$ $\leq 1,$
$\forall(i,j)\in\mathcal{D_{F}}\times\mathcal{D_{T}}$
@@
Using the predefined influence functions in Appendix B, a small fraction of influence values are computed – we call this fraction $u$. We use $u$of data from $\mathcal{D_{F}}$ and $u$of data from $\mathcal{D_{T}}$ to compute the training set for the InfluenceNetwork. As mentioned above, this training set is represented by Q1 in the illustration.
@@
To ensure our InfluenceNetwork is able to output influence values correctly, we compute the average mean squared error (MSE) between the ground truth influence values (from Appendix B) and the predicted influence values:
@@
4.4 Hyperparameter Study #1: InfluenceNetwork sizes
@@
Dataset MixInstruct Alpaca Method ICL QLoRA ICL QLoRA Metric ROUGE BGE LAJ ROUGE BGE LAJ ROUGE BGE LAJ ROUGE BGE LAJ Initial 37.87 78.92 2.98 36.36 82.55 3.02 25.79 67.82 2.56 27.29 71.57 2.62 Random 39.00 80.66 3.12 44.45 85.46 3.12 34.93 73.50 3.07 35.57 75.16 2.96 SelectIT 43.08 84.50 3.18 45.14 85.88 3.21 33.56 77.10 3.12 34.04 78.10 3.21 DistilGPT2 + SelectIT 40.21 79.37 3.05 41.60 81.75 3.08 31.68 74.86 3.04 32.30 75.75 3.14 NN-CIFT + SelectIT 43.71 81.95 3.16 46.09 86.13 3.19 34.85 77.79 3.13 34.07 78.11 3.16 LESS 42.08 83.24 3.26 45.16 84.95 3.28 35.78 76.84 3.16 35.28 76.49 3.15 DistilGPT2 + LESS 40.33 79.25 3.19 42.57 79.48 3.17 32.91 74.19 3.09 35.85 76.64 3.15 NN-CIFT + LESS 42.84 83.74 3.26 45.18 84.63 3.26 36.12 77.11 3.16 36.49 75.75 3.16 DELIFT (SE) 47.43 84.40 3.28 48.22 86.50 3.28 37.53 80.76 3.25 42.66 84.26 3.18 DistilGPT2 + DELIFT(SE) 46.74 84.36 3.23 45.50 84.06 3.29 35.99 79.38 3.20 40.15 83.89 3.09 NN-CIFT + DELIFT (SE) 47.30 82.99 3.23 46.49 84.68 3.29 37.02 80.72 3.26 42.52 84.58 3.29 DELIFT 48.46 85.77 3.35 52.79 88.04 3.37 38.36 81.13 3.36 43.43 85.05 3.56 DistilGPT2 + DELIFT 42.49 79.19 3.19 48.34 84.60 3.28 32.50 74.49 3.25 38.26 79.58 3.46 NN-CIFT + DELIFT 48.57 83.90 3.41 53.30 81.34 3.54 38.99 80.29 3.49 44.64 85.23 3.57 Full Data 58.65 88.72 3.45 65.51 92.24 3.51 35.27 77.85 3.31 39.29 78.85 3.29
@@
Table 2: Results on the Phi-3 model with $v=0.3,u=0.05$. NN-CIFT + Method indicates using NN-CIFT to estimate influence values computed from the corresponding method’s influence function. DistilGPT2 + Method indicates using the DistilGPT2 model as the language model in the corresponding method’s influence function. The average performance difference between NN-CIFT and the original influence function is merely 1.40%.
@@
Dataset MixInstruct Alpaca Method ICL QLoRA ICL QLoRA Metric ROUGE BGE LAJ ROUGE BGE LAJ ROUGE BGE LAJ ROUGE BGE LAJ Initial 28.53 74.05 2.94 34.42 78.54 3.00 24.85 72.45 2.26 34.29 80.82 3.03 Random 40.07 84.04 3.26 41.68 84.26 3.22 36.95 80.47 3.12 38.64 80.46 3.07 SelectIT 46.51 86.18 3.25 50.31 87.38 3.25 41.42 83.25 3.27 44.51 84.18 3.34 DistilGPT2 + SelectIT 41.26 80.33 3.20 44.86 84.72 3.23 39.18 80.99 2.99 41.72 81.50 3.14 NN-CIFT + SelectIT 46.48 85.86 2.28 50.87 87.43 3.26 42.07 83.67 3.27 44.99 85.13 3.37 LESS 48.21 86.19 3.34 51.24 86.07 3.37 43.34 84.19 3.38 44.73 84.04 3.32 DistilGPT2 + LESS 42.18 78.34 3.23 48.64 79.09 3.27 42.02 80.89 3.29 42.51 82.35 3.29 NN-CIFT + LESS 48.20 86.31 3.36 51.56 86.39 3.41 44.42 84.69 3.32 46.40 85.44 3.36 DELIFT (SE) 48.36 85.91 3.38 51.43 86.20 3.34 44.30 85.52 3.41 45.35 86.34 3.48 DistilGPT2 + DELIFT (SE) 47.21 84.24 3.28 49.37 84.24 3.29 43.51 85.45 3.41 44.89 79.81 3.36 NN-CIFT + DELIFT (SE) 48.59 85.01 3.39 50.53 86.10 3.33 45.49 86.27 3.44 45.75 86.45 3.47 DELIFT 51.66 88.02 3.43 55.58 91.81 3.50 46.49 87.60 3.50 49.16 87.74 3.54 DistilGPT2 + DELIFT 47.09 84.74 3.26 48.21 84.24 3.28 45.08 81.45 3.41 41.07 83.22 3.44 NN-CIFT + DELIFT 52.03 88.38 3.41 55.85 91.96 3.51 46.26 87.41 3.55 49.15 87.74 3.50 Full Data 54.43 92.55 3.40 59.47 94.12 3.58 48.53 91.21 3.63 48.29 90.82 3.66
@@
Table 3: Results on the Llama-8B model with $v=0.3,u=0.05$. NN-CIFT + Method and DistilGPT2 + Method follow the same definitions as in Table 2. The average performance difference between NN-CIFT and the original influence function is merely 1.39%.

### v4

# Neural Networks for Learnable and Scalable Influence Estimation of Instruction Fine-Tuning Data

@@ Abstract
Influence functions provide crucial insights into model training, but existing methods suffer from large computational costs and limited generalization. Particularly, recent works have proposed various metrics and algorithms to calculate the influence of data using language models, which do not scale well with large models and datasets. This is because of the expensive forward and backward passes required for computation, substantial memory requirements to store large models, and poor generalization of influence estimates to new data. In this paper, we explore the use of small neural networks – which we refer to as the InfluenceNetwork – to estimate influence values, achieving up to 99% cost reduction. Our evaluation demonstrates that influence values can be estimated with models just 0.0007% the size of full language models (we average across 1.5B-22B versions). We apply our algorithm of estimating influence values (called NN-CIFT: Neural Networks for effiCient Instruction Fine-Tuning) to the downstream task of subset selection for general instruction fine-tuning. In our study, we include four state-of-the-art influence functions and show no compromise in performance, despite large speedups, between NN-CIFT and the original influence functions. We provide an in-depth hyperparameter analyses of NN-CIFT. The code for our method can be found here: https://github.com/agarwalishika/NN-CIFT.
@@ 1 Introduction
Table 1: Approximating the computational complexity of data valuation in terms of the cost of forward passes ($F$) or backward passes ($B$) through a model. $M=|\mathcal{D_{F}}|$ and $N=|\mathcal{D_{T}}|$, a fine-tuning and target dataset respectively, we use for subset selection. See Appendix C.1 for more details. Size denotes the number of parameters of the corresponding model. Note: larger models have a higher $F$ and $B$.
In this paper, we introduce NN-CIFT: Neural Networks for effiCient Instruction Fine-Tuning and explore how to train influence functions efficiently. We improve efficiency by using compact neural networks – which we coin as the InfluenceNetwork – that are 0.0077% the size of LLMs, to estimate influence. Figure 1 outlines our methodology with a pairwise influence function (more details about pairwise influence functions in Appendix C.1).
@@
    alleviates the cost of using expensive LLMs during data valuation by using smaller and cheaper neural networks, without affecting the performance on downstream tasks (Tables 4-6);
@@
Section 2 outlines the current state of research in data valuation and data selection. Section 3 explains the problem setting. Section 4 presents the main methodology for NN-CIFT and motivating results. Finally, Section 5 reports results on the downstream task of subset selection after the data valuation stage. In our evaluation, we find that using a small LLM with the original influence functions results in degraded performance. Our hyperparameter studies are in Appendix A.1, Figure 4 and Appendix A.2, Figure 8. We also show language model performance with smaller subsets of selected fine-tuning data in Appendix B. Lastly, the SOTA influence functions are detailed in Appendix C.
@@ 3 Problem Formulation
As mentioned before, computing influence functions can be a very expensive process. There are two kinds of influence functions: pairwise and pointwise – both require forward/backward passes through language models, but the costs slightly differ. Pairwise influence functions compute the influence between every pair of points in a dataset. We study three SOTA pairwise functions, whose formulations are details in Appendix C.1. This paper also studies one pointwise influence functions that simply compute the influence of each data point individually, formally outlined in Appendix C.2. While pointwise influence functions are more efficient than pairwise, they are not as performant during subset selection Xia et al. (2024); Agarwal et al. (2025).
@@ 4 Learning Influence Estimation
$0\leq IN_{\theta}(i,j)$ $\leq 1,$ (1)
$0\leq\theta(\texttt{concat}(\texttt{emb}(i),\texttt{emb}(j)))$ $\leq 1,$ (2)
$\forall(i,j)\in\mathcal{D_{F}}\times\mathcal{D_{T}}$ (3)
@@
We note that NN-CIFT relies on an embedding model that is often larger than the actual size of the InfluenceNetwork. However, we choose not to include it in the cost for a few reasons: (1) our method does not rely on the underlying embedding model, (2) NLP pipelines generally use embedding models to store, retrieve, cluster, and/or visualize data and hence, is an offline cost.
@@
Using the predefined influence functions in Appendix C, a small fraction of influence values are computed – we call this fraction $u$. We use uu% of data from $\mathcal{D_{F}}$ and uu% of data from $\mathcal{D_{T}}$ to compute the training set for the InfluenceNetwork. As mentioned above, this training set is represented by Q1 in the illustration.
@@
To ensure our InfluenceNetwork is able to output influence values correctly, we compute the average mean squared error (MSE) between the ground truth influence values (from Appendix C) and the predicted influence values:
@@
Finally, we report the robustness of the InfluenceNetwork. In Table 2 reports the variance of the InfluenceNetwork after five runs. We assume the original influence values are given and static, and therefore, do not measure the variance of the individual influence functions. Table 2 shows low variance across each quadrant and each $u$ in the InfluenceNetwork scores. Table 3 shows that the InfluenceNetwork is invariant to the choice of the embedding model. As the MSE between the predicted influence values and ground truth influence values remains small, we posit the downstream performance of NN-CIFT will be maintained no matter the choice of embedding model.
u Q1 std Q2 std Q3 std Q4 std
0.01 0.0050 0.0047 0.0018 0.0118
0.05 0.0090 0.0091 0.0032 0.0032
0.10 0.0062 0.0058 0.0032 0.0032
0.15 0.0062 0.0056 0.0037 0.0037
0.20 0.0102 0.0096 0.0020 0.0020
0.25 0.0102 0.0096 0.0020 0.0020
0.30 0.0118 0.0112 0.0020 0.0020
0.40 0.0067 0.0053 0.0018 0.0018
Table 2: Variance of the InfluenceNetwork for varying $u$’s for DELIFT values.
Embedding Model Q1 Q2 Q3 Q4
BAAE/bge-large-en-v1.5 0.051 0.084 0.074 0.084
Qwen/Qwen3-Embedding-0.6B 0.076 0.087 0.087 0.089
intfloat/e5-mistral-7b-instruct 0.026 0.083 0.084 0.083
Snowflake/snowflake-arctic-embed-l-v2.0 0.077 0.087 0.087 0.088
Table 3: Varying embedding models and their corresponding MSE values (averaged across 5 runs) between estimated influence values and ground truth influence values, with 5% selected data. These results are for learning DELIFT influence values. As shown, NN-CIFT is invariant to the embedding model selected, and is able to effectively estimate the influence values.
@@
Dataset MixInstruct Alpaca MMLU Method ICL QLoRA ICL QLoRA ICL QLoRA Metric ROUGE BGE LAJ ROUGE BGE LAJ ROUGE BGE LAJ ROUGE BGE LAJ Accuracy Accuracy Initial 28.53 74.05 2.94 34.42 78.54 3.00 24.85 72.45 2.26 34.29 80.82 3.03 70.7 70.8 Random 40.07 84.04 3.26 41.68 84.26 3.22 36.95 80.47 3.12 38.64 80.46 3.07 71.9 71.9 SelectIT 46.51 86.18 3.25 50.31 87.38 3.25 41.42 83.25 3.27 44.51 84.18 3.34 72.7 73.0 + DistilGPT2 41.26 80.33 3.20 44.86 84.72 3.23 39.18 80.99 2.99 41.72 81.50 3.14 72.0 72.7 + NN-CIFT 46.48 85.86 2.28 50.87 87.43 3.26 42.07 83.67 3.27 44.99 85.13 3.37 74.7 72.9 LESS 48.21 86.19 3.34 51.24 86.07 3.37 43.34 84.19 3.38 44.73 84.04 3.32 75.6 76.7 + DistilGPT2 42.18 78.34 3.23 48.64 79.09 3.27 42.02 80.89 3.29 42.51 82.35 3.29 74.6 74.1 + NN-CIFT 48.20 86.31 3.36 51.56 86.39 3.41 44.42 84.69 3.32 46.40 85.44 3.36 75.0 76.5 DELIFT (SE) 48.36 85.91 3.38 51.43 86.20 3.34 44.30 85.52 3.41 45.35 86.34 3.48 78.9 79.8 + DistilGPT2 47.21 84.24 3.28 49.37 84.24 3.29 43.51 85.45 3.41 44.89 79.81 3.36 75.4 76.1 + NN-CIFT 48.59 85.01 3.39 50.53 86.10 3.33 45.49 86.27 3.44 45.75 86.45 3.47 78.5 79.8 DELIFT 51.66 88.02 3.43 55.58 91.81 3.50 46.49 87.60 3.50 49.16 87.74 3.54 81.5 83.1 + DistilGPT2 47.09 84.74 3.26 48.21 84.24 3.28 45.08 81.45 3.41 41.07 83.22 3.44 77.1 78.5 + NN-CIFT 52.03 88.38 3.41 55.85 91.96 3.51 46.26 87.41 3.55 49.15 87.74 3.50 82.0 83.6 Full Data 54.43 92.55 3.40 59.47 94.12 3.58 48.53 91.21 3.63 48.29 90.82 3.66 80.5 81.6
@@
Table 4: Results on the Llama-8B model with $v=0.3,u=0.05$. “+ NN-CIFT” indicates using NN-CIFT to estimate influence values computed from the corresponding method’s influence function. “+ DistilGPT2” indicates using the DistilGPT2 model as the language model in the corresponding method’s influence function. The average performance difference between NN-CIFT and the original influence function is merely 0.13%.
@@
Dataset MixInstruct Alpaca MMLU Method ICL QLoRA ICL QLoRA ICL QLoRA Metric ROUGE BGE LAJ ROUGE BGE LAJ ROUGE BGE LAJ ROUGE BGE LAJ Accuracy Accuracy Initial 16.19 61.32 2.06 19.31 64.27 2.09 24.53 71.42 2.48 24.80 71.79 2.61 69.4 70.9 Random 28.33 72.41 2.37 29.93 75.78 2.50 26.54 72.71 2.66 28.10 73.00 2.78 68.9 71.9 SelectIT 40.64 72.63 2.21 44.85 75.72 2.83 31.86 76.29 2.73 32.56 78.17 2.77 71.7 71.6 + DistilGPT2 40.33 71.49 2.11 43.26 74.14 2.38 29.65 75.32 2.62 31.66 74.20 2.67 69.9 70.8 + NN-CIFT 41.59 71.36 2.26 45.87 74.22 2.58 32.05 76.82 2.78 32.90 77.12 2.80 71.1 72.0 LESS 45.33 78.68 3.03 46.03 81.04 3.05 38.43 78.83 2.98 41.68 81.83 3.09 74.0 75.6 + DistilGPT2 43.15 74.69 2.42 42.87 75.80 2.46 34.76 75.26 2.86 37.66 78.57 3.06 66.7 69.3 + NN-CIFT 46.32 78.84 3.05 47.84 80.48 3.03 39.59 78.75 2.89 42.06 81.06 3.08 74.1 74.9 DELIFT (SE) 46.68 81.01 3.12 48.42 83.67 3.15 42.52 81.21 3.12 43.83 84.35 3.26 74.0 75.1 + DistilGPT2 45.89 79.77 3.07 46.07 80.40 3.10 41.15 79.42 2.96 42.32 82.63 3.02 75.0 75.4 + NN-CIFT 46.81 81.23 3.14 48.64 82.76 3.16 42.72 80.86 3.16 42.75 84.52 3.27 75.2 75.5 DELIFT 48.85 83.89 3.25 50.90 85.64 3.27 44.74 83.60 3.28 46.33 87.87 3.46 77.4 79.4 + DistilGPT2 43.69 77.83 3.04 45.61 79.76 3.06 40.25 77.10 2.96 43.37 79.88 2.96 74.6 75.4 + NN-CIFT 48.97 83.57 3.25 49.71 86.45 3.29 45.78 85.49 3.29 48.69 87.14 3.48 77.5 79.3 Full Data 49.31 86.25 3.48 52.55 89.58 3.51 49.31 89.39 3.45 50.95 90.23 3.66 73.7 74.6
@@
Table 5: Results on the Qwen2.5 with $v=0.3,u=0.05$. “+ NN-CIFT” indicates using NN-CIFT to estimate influence values computed from the corresponding method’s influence function. “+ DistilGPT2” indicates using the DistilGPT2 model as the language model in the corresponding method’s influence function. The average performance difference between NN-CIFT and the original influence function is merely 0.24%.
Dataset MixInstruct Alpaca MMLU Method ICL QLoRA ICL QLoRA ICL QLoRA Metric ROUGE BGE LAJ ROUGE BGE LAJ ROUGE BGE LAJ ROUGE BGE LAJ Accuracy Accuracy Initial 32.31 74.27 2.18 35.16 78.13 2.21 37.44 78.65 2.41 37.81 78.69 2.43 71.9 72.3 Random 35.83 80.62 2.98 37.13 81.57 3.01 40.44 82.75 2.48 40.19 81.15 2.46 71.3 72.3 SelectIT 40.64 84.99 3.17 45.97 86.31 3.20 42.72 82.00 2.54 43.06 83.59 2.67 73.6 75.8 + DistilGPT2 39.84 80.48 3.02 41.81 81.21 3.03 40.74 81.44 2.37 40.36 81.85 2.53 73.5 73.8 + NN-CIFT 40.33 84.95 3.14 46.85 86.06 3.19 41.52 82.18 2.57 43.45 82.86 2.63 73.6 75.4 LESS 48.33 85.78 3.30 49.93 87.42 3.37 44.14 85.80 3.04 47.13 87.94 3.05 75.4 76.6 + DistilGPT2 45.22 82.80 3.23 44.53 84.56 3.31 40.92 80.23 2.68 42.65 84.29 2.96 74.5 74.4 + NN-CIFT 48.75 86.15 3.33 51.93 86.57 3.41 44.43 85.79 3.07 47.90 88.01 3.04 75.2 77.3 DELIFT (SE) 48.84 88.10 3.42 48.85 88.16 3.44 45.13 87.34 3.35 48.47 89.19 3.42 76.7 77.1 + DistilGPT2 46.25 86.22 3.39 46.29 87.67 3.41 44.84 86.00 3.25 46.86 86.93 3.38 75.7 77.7 + NN-CIFT 48.81 88.77 3.41 49.14 88.33 3.44 46.01 86.43 3.36 47.99 88.03 3.41 76.8 77.4 DELIFT 53.70 91.52 3.54 54.69 92.42 3.56 50.36 89.53 3.40 53.65 91.79 3.53 78.2 80.4 + DistilGPT2 46.55 88.81 3.41 47.78 89.34 3.46 45.31 83.69 3.23 44.78 85.08 3.31 75.3 76.9 + NN-CIFT 53.53 90.82 3.52 56.65 91.42 3.55 50.67 88.97 3.43 52.48 92.03 3.53 78.4 80.9 Full Data 52.30 91.32 3.60 54.83 91.57 3.69 50.61 90.77 3.41 54.51 90.70 3.58 79.1 81.2
Table 6: Results on the Mistral model with $v=0.3,u=0.05$. “+ NN-CIFT” indicates using NN-CIFT to estimate influence values computed from the corresponding method’s influence function. “+ DistilGPT2” indicates using the DistilGPT2 model as the language model in the corresponding method’s influence function. The average performance difference between NN-CIFT and the original influence function is merely 0.12%.
4.4 Interpretive Analysis
Figure 3: Distribution of influence values across each of the methods on the Alpaca dataset. To clarify, "D_F" is $\mathcal{D_{F}}$ and "D_T" is $\mathcal{D_{T}}$. The x-axis spans the 15,000 examples from $\mathcal{D_{F}}$ and the y-axis spans the 5,000 samples from $\mathcal{D_{T}}$. SelectIT only has an x-axis.
We posit two reasons that contribute to the InfluenceNetwork’s success: (1) most data points have the same influence, and (2) influence estimation is a lossy task.
Firstly, we visualize the distribution of the influence values in Figure 3. As shown, the most "different" influence values (the stark dark/light colors indicating low/high influence) are sparse. This indicates that the neural network simply has to learn to estimate the extremes, and can achieve good performance for the rest of the values. Second, we notice that influence estimation is an incredibly lossy task. It involves compressing high-dimensional text representations to a singular, scalar value. Furthermore, the scalar values are constrained to a small range (see Equation 1), reducing the margin of error. Putting these two reasons together, we can see why a neural network can replace a language model during influence estimation without any significant effect on performance.